# Motives of nullcones of quiver representations

### v2

@@ 5. Wall-crossing type formula
We define a new twisted formal power series ring $R_{\mathbb{L},*}\llbracket(Q_{\mathbb{Z}})_{0}\rrbracket$ for $Q_{\mathbb{Z}}$ with topological $R$-basis elements $u^{\bf e}$ for ${\bf e}$ a dimension vector for $Q_{\mathbb{Z}}$ with finite support, and multiplication given by
$u^{\mathbf{e}}\ast u^{\mathbf{e}^{\prime}}=\mathbb{L}^{-\sum_{k\geq 0}\langle\mathbf{e},{\tau}^{k}\mathbf{e}^{\prime}\rangle_{\mathbb{Q}_{\mathbb{Z}}}-\sum_{k>0}\langle\mathbf{e}^{\prime},{\tau}^{k}\mathbf{e}\rangle_{\mathbb{Q}_{\mathbb{Z}}}}\cdot u^{\mathbf{e}+\mathbf{e}^{\prime}}~{}.$
We have a specialization map of $R$-modules $R_{\mathbb{L},*}\llbracket(Q_{\mathbb{Z}})_{0}\rrbracket\rightarrow R_{\mathbb{L}}\llbracket Q_{0}\rrbracket$ by mapping $u^{\bf e}$ to $t^{|{\bf e}|}$ (in general not compatible with the multiplications).
This definition now allows us to directly conclude the following product formula using an ascending ordered product:
@@
The nilpotent motivic generating function of Proposition 5.4 equals the specialization of the following ordered product in $R_{\mathbb{L},*}\llbracket(Q_{\mathbb{Z}})_{0}\rrbracket$:
$\prod^{\rightarrow}_{\mu\in[0,1[\cap\mathbb{Q}}\Biggl{(}\sum\limits_{\begin{subarray}{c}\mathbf{e}\in\mathbb{N}(Q_{\mathbb{Z}})_{0}\\ \operatorname{supp}(\mathbf{e})\text{ finite}\end{subarray}}\mathbb{L}^{-\sum_{k>0}\langle\mathbf{e},{\tau}^{k}\mathbf{e}\rangle}\frac{[R_{\mathbf{e}}^{\theta-sst}(Q_{\mathbb{Z}})]}{[G_{\mathbf{e}}]}\cdot u^{\mathbf{e}}\Biggr{)}~{}.$

### v3

@@ 5 Wall-crossing type formula
We define a new twisted formal power series ring $\mathbb{M}_{\mathbb{L},*}\llbracket(Q_{\mathbb{Z}})_{0}\rrbracket$ for $Q_{\mathbb{Z}}$ with topological $\mathbb{M}$-basis elements $u^{\bf e}$ for $\bf e$} a dimension vector for $Q_{\mathbb{Z}}$ with finite support, and multiplication given by
$u^{\mathbf{e}}\ast u^{\mathbf{e}^{\prime}}=\mathbb{L}^{-\sum_{k\geq 0}\langle\mathbf{e},{\tau}^{k}\mathbf{e}^{\prime}\rangle_{\mathbb{Q}_{\mathbb{Z}}}-\sum_{k>0}\langle\mathbf{e}^{\prime},{\tau}^{k}\mathbf{e}\rangle_{\mathbb{Q}_{\mathbb{Z}}}}\cdot u^{\mathbf{e}+\mathbf{e}^{\prime}}.$
We have a specialization map of $\mathbb{M}$-modules $\mathbb{M}_{\mathbb{L},*}\llbracket(Q_{\mathbb{Z}})_{0}\rrbracket\rightarrow\mathbb{M}_{\mathbb{L}}\llbracket Q_{0}\rrbracket$ by mapping $u^{\bf e}$ to $t^{|{\bf e}|}$ (in general not compatible with the multiplications).
This definition now allows us to interpret Proposition 5.4 as a product formula using an ascending ordered product over slopes. Namely, given series $F_{\mu}\in\mathbb{M}_{\mathbb{L},*}\llbracket(Q_{\mathbb{Z}})_{0}\rrbracket$ without constant term for $\mu\in[0,1]\cap\mathbb{Q}$ as below, we define
$\prod_{\mu\in[0,1]\cap\mathbb{Q}}^{\rightarrow}(1+F_{\mu})=\sum_{s\geq 0}\,\,\sum_{0\leq\mu_{1}<\dots<\mu_{s}<1}\,\,\prod_{k=1}^{s}F_{\mu_{k}}.$
Then we find by a direct calculation the following.
@@
The nilpotent motivic generating function of Proposition 5.4 equals the specialization of the following ordered product in $\mathbb{M}_{\mathbb{L},*}\llbracket(Q_{\mathbb{Z}})_{0}\rrbracket$:
$\prod^{\rightarrow}_{\mu\in[0,1]\cap\mathbb{Q}}\Biggl(1+\sum\limits_{\begin{subarray}{c}\sigma({\bf e})=\mu\\ \operatorname{supp}(\mathbf{e})\ \mathrm{finite}\end{subarray}}\mathbb{L}^{-\sum_{k>0}\langle\mathbf{e},{\tau}^{k}\mathbf{e}\rangle}\frac{\big[R_{\mathbf{e}}^{\theta\text{\rm-sst}}(Q_{\mathbb{Z}})\big]}{[G_{\mathbf{e}}]}\cdot u^{\mathbf{e}}\Biggr).$
Acknowledgments
The authors would like to thank anonymous referees for suggestions on improving the presentation and for pointing out [1, Lemma 2.6]. The second named author would like to thank S. Mozgovoy for stimulating discussions. The work of both authors is supported by the DFG GRK 2240 “Algebro-Geometric Methods in Algebra, Arithmetic and Topology”.